OpenStudy (anonymous):

5(x + 10) + x 6x + 15 5x + 15 6x + 50 4x – 50

OpenStudy (anonymous):

5(x+10) = 5x + 50 (distribute) 5x + 50 + x = 6x +50 so the 3rd one

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

i have another 1 hang on

OpenStudy (anonymous):

np

OpenStudy (anonymous):

–4n + 7 + 2n = 1 1 3 -3 4

OpenStudy (anonymous):

You need to put n together first

OpenStudy (anonymous):

so -4n +2n =...

OpenStudy (anonymous):

-2n

OpenStudy (anonymous):

then move numbers to the otherside, -2n = 1-7

OpenStudy (anonymous):

So...?

OpenStudy (anonymous):

-2n = -6, n =?

OpenStudy (anonymous):

divide both sides by -2, so you will get n= -6/-2

OpenStudy (anonymous):

so its 3?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

4(y – 4) = 8 –2 2 4 6

OpenStudy (anonymous):

ok 1 more then thats it

OpenStudy (anonymous):

i just posted it

OpenStudy (anonymous):

did you try it ?

OpenStudy (anonymous):

Try what

OpenStudy (anonymous):

solve for y ?

OpenStudy (anonymous):

unless you stuck

OpenStudy (anonymous):

yea im stuck

OpenStudy (anonymous):

FIrst, distribute out, 4y -16 =8 4y = 24, y =6

OpenStudy (anonymous):

Wow. Now. I. Feel. Stupid. LOL--thanks so much!

OpenStudy (anonymous):

thats it!!!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

im gonna close this Q now--thanks again!

OpenStudy (anonymous):

np